# Identifying Important Group of Pixels using Interactions

### v1

@@ D Additional results of visualization
E Class-descriminative localization
The proposed method was originally designed to identify important pixels to explain the model prediction. Here, we generalize MoXI (for pixel deletion) to visualize such pixels for a given target class, which is used in Fig. 5.
To this end, we consider reward function switching as follows. Let $x,y_{\mathrm{t}},y_{f(x)}$ be the input image, the target label, and the predicted label, respectively. If $y_{\mathrm{t}}=y_{f(x)}$, we simply use a reward function $f(x)=\log\frac{P(y_{\mathrm{t}}\,|\,x)}{1-P(y_{\mathrm{t}}\,|\,x)}$. Otherwise, we use $f(x)=\log\frac{P(y_{f(x)}\,|\,x)}{1-Py_{f(x)}\,|\,x)}-\frac{P(y_{\mathrm{t}}\,|\,x)}{1-P(y_{\mathrm{t}}\,|\,x)}$, which helps us identify patches with positive effect on the confidence score on class $y_{f(x)}$ and negative effect on class $y_{\mathrm{t}}$. The image patches removed in the former case are collected as important patches for class $y_{\mathrm{t}}$.
F Patch perturbations
@@ F.1 Common corruptions
We implemented 19 types of common corruptions using the imagecorruptions module with severity 5.555 https://github.com/hendrycks/robustness. Figures 11 and 12 showcase the deletion curves with different corruptions for ViT-T and DeiT-T, respectively.
The results demonstrate that our method gives a sharper decrease at the early stage of deletion curves than others, as in Sec. 5.3.
Figure 11: Deletion curves with ViT-T. The curves show the accuracy decay when replacing clean patches with corrupted ones according to their contributions calculated by each method. The horizontal axis presents the replacement rate.
Figure 12: Deletion curves with DeiT-T. The curves show the accuracy decay when replacing clean patches with corrupted ones according to their contributions calculated by each method. The horizontal axis presents the replacement rate.
F.2 Adversarial perturbations
Besides common corruptions, we also investigated the case with adversarial perturbations [12, 17, 19], which are small but malicious perturbations that can largely change the model’s output. We conducted the same experiment given in Sec. F.1 but with adversarial perturbations instead of common corruptions. To obtain adversarial perturbations, we adopted $L_{2}$-untargeted PGD with $\epsilon=1.0$ and stepsize $\alpha=0.2$. Figure 14(a) and 14(a) present the deletion curves for ViT-T and Deit-T, respectively. The results show that the attention rollout method gives a slightly sharper decrease than MoXI. This differs from the results for common corruptions. We suspect that adversarial perturbations mostly lie in the patches that are suggested as important by attention rollout.
@@ F.2 Adversarial perturbations
To confirm this, we measured the magnitude of adversarial perturbations on each image patch. Specifically, the magnitude is measured by the L2 norm. Figure 14(b) shows the magnitude of the perturbations of each patch. The patches are ordered as in the deletion curves in Fig. 14(a). The results indicate that the importance of image patches identified by attention rollout is well aligned with the amount of perturbations on them. On the other hand, image patches identified by MoXI contain a larger amount of perturbations at the early and late stages than those at the middle stage. This may be because the attention rollout reflects the internal computation process of the features directly when measuring the contributions of image patches, while adversarial perturbations are designed to hack this process. On the other hand, MoXI treats a Vision Transformer as a black-box model and is unaware of the internal process.
Figure 13: (a) Deletion curves with ViT-T. The curves show the accuracy decay when replacing clean patches with adversarial ones according to their contributions calculated by each method. The horizontal axis presents the replacement rate. (b) The amount of adversarial perturbations.
Figure 13: (a) Deletion curves with ViT-T. The curves show the accuracy decay when replacing clean patches with adversarial ones according to their contributions calculated by each method. The horizontal axis presents the replacement rate. (b) The amount of adversarial perturbations.
Figure 14: (a) Deletion curves with DeiT-T. The curves show the accuracy decay when replacing clean patches with adversarial ones according to their contributions calculated by each method. The horizontal axis presents the replacement rate. (b) The amount of adversarial perturbations.
G More results in the stability of explanations.
@@
Figure 15: (Top) Insertion curves. (Bottom) Deletion curves. The curves illustrate the accuracy growth when inserting (deleting) image patches according to the contributions computed by each method. The horizontal axis presents the insertion rate. (a) Grad-CAM results, (b) Attention Rollout results, (c) Shapley Value results, (d) MoXI results.

### v2

@@ E Additional results of visualization
F Class-descriminative localization
The proposed method was originally designed to identify important pixels to explain the model prediction. Here, we generalize MoXI (for pixel deletion) to visualize such pixels for a given target class, which is used in Fig. 5. To this end, we consider reward function switching as follows. Let $x,y_{\mathrm{t}},y_{f(x)}$ be the input image, the target label, and the predicted label, respectively. If $y_{\mathrm{t}}=y_{f(x)}$, we simply use a reward function $f(x)=\log\frac{P(y_{\mathrm{t}}\,|\,x)}{1-P(y_{\mathrm{t}}\,|\,x)}$. Otherwise, we use $f(x)=\log\frac{P(y_{f(x)}\,|\,x)}{1-Py_{f(x)}\,|\,x)}-\frac{P(y_{\mathrm{t}}\,|\,x)}{1-P(y_{\mathrm{t}}\,|\,x)}$, which helps us identify patches with positive effect on the confidence score on class $y_{f(x)}$ and negative effect on class $y_{\mathrm{t}}$. The image patches removed in the former case are collected as important patches for class $y_{\mathrm{t}}$.
G Patch perturbations
@@ G.1 Common corruptions
We implemented 19 types of common corruptions using the imagecorruptions module with severity 5.666 https://github.com/hendrycks/robustness. Figures 13 and 14 showcase the deletion curves with different corruptions for ViT-T and DeiT-T, respectively. The results demonstrate that our method gives a sharper decrease at the early stage of deletion curves than others, as in Sec. 5.3.
Figure 13: Deletion curves by image corruptions instead of masking with ViT-T. The curves illustrate the change in accuracy along with the increase in the number of corrupted image patches. The patches are corrupted from the highly contributing ones determined by each method.
Figure 14: Deletion curves by image corruptions instead of masking with DeiT-T. The curves illustrate the change in accuracy along with the increase in the number of corrupted image patches. The patches are corrupted from the highly contributing ones determined by each method.
G.2 Adversarial perturbations
@@
Besides common corruptions, we also investigated the case with adversarial perturbations [12, 17, 19], which are small but malicious perturbations that can largely change the model’s output. We conducted the same experiment given in Sec. G.1 but with adversarial perturbations instead of common corruptions. To obtain adversarial perturbations, we adopted $L_{2}$-untargeted PGD with $\epsilon=1.0$ and stepsize $\alpha=0.2$. Figure 16(a) and 16(a) present the deletion curves for ViT-T and Deit-T, respectively. The results show that the attention rollout method gives a slightly sharper decrease than MoXI. This differs from the results for common corruptions. We suspect that adversarial perturbations mostly lie in the patches that are suggested as important by attention rollout. To confirm this, we measured the magnitude of adversarial perturbations on each image patch. Specifically, the magnitude is measured by the L2 norm. Figure 16(b) shows the magnitude of the perturbations of each patch. The patches are ordered as in the deletion curves in Fig. 16(a). The results indicate that the importance of image patches identified by attention rollout is well aligned with the amount of perturbations on them. On the other hand, image patches identified by MoXI contain a larger amount of perturbations at the early and late stages than those at the middle stage. This may be because the attention rollout reflects the internal computation process of the features directly when measuring the contributions of image patches, while adversarial perturbations are designed to hack this process. On the other hand, MoXI treats a Vision Transformer as a black-box model and is unaware of the internal process.
Figure 15: (a) Deletion curves by adversarial perturbations instead of masking with ViT-T. The curves illustrate the change in accuracy along with the increase in the number of perturbed image patches. The patches are perturbed from the highly contributing ones determined by each method. (b) The amount of adversarial perturbations.
Figure 16: (a) Deletion curves by adversarial perturbations instead of masking with DeiT-T. The curves illustrate the change in accuracy along with the increase in the number of perturbed image patches. The patches are perturbed from the highly contributing ones determined by each method. (b) The amount of adversarial perturbations.
H More results in the stability of explanations.
@@
Figure 17: (Top) Insertion curves. (Bottom) Deletion curves. The curves illustrate the change in accuracy along with the increase (decrease) in the number of unmasked (masked) image patches. Each curve represents the results from the pretrained models with $10$, $20$, $100$, and $1000$ classes, respectively. (a) Grad-CAM results, (b) Attention Rollout results, (c) Shapley Value results, (d) MoXI results.